### v2

@@ 3 Related Work
CCNet Wenzek et al. (2020) (ccnet) is an automated preprocessing pipeline that aims to extract high-quality monolingual language datasets from the Common Crawl corpus of Web content. CCNet performs dataset sharding, deduplication, language identification, and quality filtering. CCNet’s deduplication technique begins by preprocessing text so that all characters are lowercase and eliminates special Unicode characters. CCNet then splits documents by newline characters, computes SHA1 hashes for each resulting unit of text, and compares SHA1 hashes between units of text to identify duplicates. CCNet can optionally compare to a single document, a subset of documents, or all documents.
Dolma Soldaini et al. (2024) deduplicates paragraphs by exact matching using a Bloom filter (dolma). An alternative procedure (dolma-ngram) splits paragraphs into n-grams and checks whether n-grams are duplicated by querying against a Bloom filter. If the proportion of duplicated n-grams in a paragraph exceeds a threshold, $T\in$ [0, 1], then that paragraph is marked as a duplicate. By default, $T$ = 1.0 in their configuration.
DataComp-LM Li et al. (2024) (dclm) deduplicates items at the document and paragraph levels simultaneously by using a Big Friendly Filter, an extension of the Bloom filter. They split each document into paragraphs on newline characters and tokenize paragraphs with the UniSeg tokenizer Ye et al. (2023). For each paragraph, they query each n-gram against their Bloom filter and remove duplicated n-grams. If the proportion of duplicated n-grams in a paragraph exceeds a user-defined threshold, $T\in$ [0, 1], the entire paragraph is removed from the document. If the proportion of duplicated n-grams in the entire document exceeds the user-defined threshold, then the document is removed from the dataset. They employ the same threshold parameters for paragraph-level and document-level deduplication.
4 The LSHBloom Method
4.1 Overview
We propose a memory-efficient alternative to MinHashLSH for document deduplication. In particular, rather than employ the traditional prefix-tree index used to find exact matches between the MinHash signatures in each band, we develop a novel Bloom filter index structure for MinHashLSH. We instantiate a series of Bloom filters, one corresponding to each band of the MinHash signature matrix. By inserting several rows of MinHash signatures into a fixed number of Bloom filters of fixed size, we achieve a large reduction in space usage by the index and a significant speedup in searching for duplicates. The size of each Bloom filter and the number of hash functions utilized are determined according to the optimal filter size as computed by the method of Bender et al. (2022).
4.2 Insertion
@@
When we wish to insert a document into the index, we first compute the MinHash signature for a document, which is a vector of integers (each produced by a universal hash function). Then, we group the rows into $b$ bands of size $r$, as dictated by the MinHashLSH algorithm. It is desirable to hash this list of integers into a single value that we can readily insert into a Bloom filter in a manner that minimizes collisions. To do so, we take advantage of the fact that each MinHash signature value is computed using a universal hash function (in our case, SHA1 3rd & Jones (2001)), with the property that the probability of a hash collision is $1/N$ where $N$ is the range of possible hash values. Then, we can hash a vector $\bar{x}$ of $r$ integers as follows Carter & Wegman (1977):
@@
This approach enables us to map each band to a single integer value with known collision probability. To insert a document into the index, we simply insert each of the $b$ final integer values into each of the $b$ Bloom filters corresponding to the $b$ bands in the signature matrix.
4.3 Querying
Querying a document against the index is a similar process to insertion. We compute the document’s MinHash signature, group the rows into bands, and then hash each band as described in Section 4.2. If there is a collision for any band of signatures, we mark that document as a duplicate. This approach mimics the original MinHashLSH procedure and gives the same probabilistic guarantees with respect to the Jaccard similarity of candidate pairs, with a slight increase in false positive probability due to our use of Bloom filters. However, the false positive probability is configurable and can be set to be arbitrarily low (in our experiments, we set the effective false positive rate as low as 1e-5).
4.4 Error Rate
@@
[…]
@@
The relationship between the false positive rate of the Bloom filter and the overall error rate of LSHBloom can be understood as follows. Leskovec et al. (2014) determine that the false positive and false negative rate of MinHashLSH can be given by the following integrals, where $T$ is the Jaccard similarity threshold, $b$ is the number of bands in the MinHash signature matrix, and $r$ is the number of rows in each band:
$FP=\int_{0}^{T}1-(1-t^{r})^{b}dt$ (1)
$FN=\int_{T}^{1}1-(1-(1-t^{r})^{b})dt$ (2)
@@
Zhu et al. (2016) demonstrate how to tune the values of $b$ and $r$ to minimize these errors in MinhashLSH. We follow this approach in our work.
@@
A Bloom filter will not yield false negatives, so the overall false negative rate is unaffected in LSHBloom. If MinHashLSH yields a false positive (by “mistakenly” yielding two identical bands), we get a false positive result. If MinHashLSH does not yield a false positive, the bloom filter index may still yield a false positive result with probability $p_{effective}$. Therefore, the overall false positive rate for LSHBloom is given by:
$FP_{total}=FP+(1-FP)*p_{effective}$
Note that $p_{effective}$ can be set arbitrarily small in practice so that this overhead is negligible.
4.5 Advantages
@@
We seek to measure duplication based on the Jaccard similarity between set-theoretic representations of documents as bags of n-grams. Since this representation stems from the k-shingling of documents and many practical scenarios may admit an extremely large universe of such k-shingles (e.g., if we allow our set of documents to include a variety of encodings, latex, tablature, figures, etc.), practitioners may want to reflect this additional complexity by increasing the range of MinHash hash values or increasing the accuracy of MinHash estimates of the Jaccard similarity. This can be done by increasing the number of permutations applied to prevent as much information loss as possible via compression of the text representation. A scalable deduplication algorithm should be amenable to such considerations. LSHBloom allows for the space-efficient representation of extremely large deduplication indices. This is due to the Bloom filter’s ability to determine the set membership properties of objects of arbitrary size with the same number of bits. If we expect $n$ documents and desire a false positive probability of $p$, then each Bloom filter requires exactly $m=-n*log(p)/(log(2))^{2}$ bits Bender et al. (2022).
@@
As an example, say we use a Jaccard similarity threshold $T$ of 0.8, and 128 random permutations for MinHashing. For these parameter settings, MinHashLSH creates nine bands for the MinHash signature matrix, meaning that we require nine Bloom filters for our index. For these settings, setting a conservative $p$effective = 1e-10 for a dataset of 10B documents would only require 590 GB of space. Meanwhile, traditional MinHashLSH may require 46 TB to store the equivalent LSHIndex—exemplifying the scalability of our technique.
@@
5 Evaluation
Broadly our evaluation considers two concerns important for large-scale deduplication: correctness and performance. As expressed above, both the correctness and performance of these methods are critical for their application to extremely large datasets. For correctness, we compare LSHBloom against state-of-the-art deduplication methods when applied to a dataset of 25K PDFs containing labeled duplicates. Without correctness, the system may either under-identify duplicates, resulting in decreased downstream inference quality; or over-identify duplicates, unnecessarily decreasing the volume of the training dataset. For performance, we consider the storage utilization and runtime of each method on a larger collection of PDFs, peS2o Soldaini & Lo (2023). Without performance, it may not be possible to process large datasets on available hardware resources.
We conduct a comprehensive analysis of various deduplication methods. Specifically, as each method includes a set of parameters whose values affect both their correctness and resource usage, we evaluate each using a range of parameter settings. Understanding these relationships is crucial to making the best use of available resources.
@@ 5.1.2 Baselines
We compare LSHBloom with state-of-the-art deduplication methods widely used in large-scale LLM training: MinHashLSH, Dolma, Dolma-ngram, CCNet, and DataComp-LM. We normalized all implementations to Python-only implementations to avoid performance differences due to implementation language. This also enabled us to use the same Bloom filter implementation for each technique. Below, we describe any changes or configurations to the methods.
@@
MinhashLSH Broder (1997) (lsh): We configure the traditional MinhashLSH algorithm as described in Section 2.3. DOLMA: We apply both standard dolma and dolma-ngram methods to document-level deduplication. Specifically, for dolma we determine whether a document is duplicated by identifying duplicated paragraphs and comparing the percentage of document text duplicated to the overlap threshold, $T$. For dolma-ngram, we split each document into a set of n-grams with a given size, determine how many n-grams are duplicated according to the Bloom filter, and compare the percentage of n-grams that are duplicated within that document to the overlap threshold, $T$. Our implementation aims to faithfully extend Dolma’s two primary deduplication methods to document-level deduplication. CCNet originally performs deduplication at the paragraph level; however, here, we apply it at the document level by measuring the proportion of duplicated paragraphs in any given document against some tolerance threshold, $T\in$ [0, 1]. For a fair comparison, we compare documents to all other documents—just as MinhashLSH does. When evaluating memory usage, we apply all CCNet optimizations, such as the use of memory-efficient hash sets. Additionally, we augment CCNet with multiprocessing where appropriate (e.g., when computing hashes over each document) for a more appropriate comparison during our scaling tests. DataComp-LM: For our evaluation, we compare only with their document-level deduplication procedure.
@@
When considering the correctness of techniques that operate on n-grams and use Bloom filters, we must give an expected n-gram count in the corpus; choosing this count accurately is crucial to make a fair comparison. Because counting the number of n-grams of a given size in a corpus is far more computationally expensive than counting the number of documents (as is required for our technique), we compute an estimate of the n-gram count in a corpus. To do so we randomly sample $N$ = 1000 documents from the corpus, compute the mean n-gram count from our sample, and multiply the mean n-gram count by the number of total documents in that corpus to obtain an estimate of the number of n-grams in the entire corpus. We parameterize Bloom filters for n-gram techniques using this estimate.
5.1.3 Metrics
@@
Correctness: We require methods that can distinguish meaningfully between duplicate and non-duplicate documents. We want these methods to yield a few false positives (to retrain the maximal number of training examples), few false negatives to reduce the prevalence of duplication in our dataset, and to provide a reasonable combination of accuracy and performance. Particularly for large-scale deep learning, retaining the maximal number of training examples is highly desirable—therefore, we want few false positives. Similarly, allowing a large number of duplicates into the training corpus is unacceptable as this affects the model’s training efficiency and may increase the likelihood that our model develops degenerate behaviors such as memorization. To account for these considerations, we evaluate each deduplication technique based on its ability to mitigate both false positives and false negatives using the F1 score:
@@
5.1.4 Constructing the correctness evaluation dataset
@@
To permit the rigorous evaluation of de-duplication methods, we created an evaluation dataset By acquiring 24,000 scientific articles (4156 scraped from HTML and 20,844 parsed from PDF) from a variety of sources, spanning a diverse set of disciplines (mathematics, physics, chemistry, biology, economics, engineering, and medicine) as identified through keywords or classification algorithms based on metadata. We then duplicated 1000 of those documents for a total of 25,000. We then translated each PDF into a sequence of tokens by applying a text extraction or OCR tool—one of PyMupDF Smith (2024), Nougat Blecher et al. (2023), Marker Paruchuri (2024), Grobid Lopez (2009), or Tesseract Smith (2007), each used with roughly the same frequency—for a total of 24,956 HTML or parsed PDF documents (as 44 failed to parse). Finally, we transformed 419 of these documents by removing a randomly selected subset of its tokens
5.1.5 Choice of values for Tuning Parameters
[…]

### v3

@@ 3. Related Work
CCNet (wenzek2020ccnet) (ccnet) is an automated preprocessing pipeline that aims to extract high-quality monolingual language datasets from the Common Crawl corpus of Web content. CCNet performs dataset sharding, deduplication, language identification, and quality filtering. CCNet’s deduplication technique begins by preprocessing text so that all characters are lowercase and eliminates special Unicode characters. CCNet then splits documents by newline characters, computes SHA1 hashes for each resulting unit of text, and compares SHA1 hashes between units of text to identify duplicates. CCNet can optionally compare to a single document, a subset of documents, or all documents.
Dolma (soldaini_dolma_2024) deduplicates paragraphs by exact matching using a Bloom filter (dolma). An alternative procedure (dolma-ngram) splits paragraphs into n-grams and checks whether n-grams are duplicated by querying against a Bloom filter. If the proportion of duplicated n-grams in a paragraph exceeds a threshold, $T\in$ [0, 1], then that paragraph is marked as a duplicate. By default, $T$ = 1.0 in their configuration.
DataComp-LM (li2024datacomp) (dclm) deduplicates items at the document and paragraph levels simultaneously by using a Big Friendly Filter, an extension of the Bloom filter. They split each document into paragraphs on newline characters and tokenize paragraphs with the UniSeg tokenizer (ye2023uniseg). For each paragraph, they query each n-gram against their Bloom filter and remove duplicated n-grams. If the proportion of duplicated n-grams in a paragraph exceeds a user-defined threshold, $T\in$ [0, 1], the entire paragraph is removed from the document. If the proportion of duplicated n-grams in the entire document exceeds the user-defined threshold, then the document is removed from the dataset. They employ the same threshold parameters for paragraph-level and document-level deduplication.
A table comparing various methods for data deduplication in contemporary LLM pre-training pipelines can be found in Appendix A.
4. The LSHBloom Method
4.1. Overview
We propose a memory-efficient alternative to MinHashLSH for document deduplication. In particular, rather than employ the traditional prefix-tree index used to find exact matches between the MinHash signatures in each band, we develop a novel Bloom filter index structure for MinHashLSH. We instantiate a series of Bloom filters, one corresponding to each band of the MinHash signature matrix. By inserting several rows of MinHash signatures into a fixed number of Bloom filters of fixed size, we achieve a large reduction in space usage by the index and a significant speedup in searching for duplicates. The size of each Bloom filter and the number of hash functions utilized are determined according to the optimal filter size as computed by the method of bender2022optimal.
4.2. Insertion
@@
When we wish to insert a document into the index, we first compute the MinHash signature for a document, which is a vector of integers (each produced by a universal hash function). Then, we group the rows into $b$ bands of size $r$, as dictated by the MinHashLSH algorithm. It is desirable to hash this list of integers into a single value that we can readily insert into a Bloom filter in a manner that minimizes collisions. To do so, we take advantage of the fact that each MinHash signature value ($h_{i}(x_{i})$) is computed using a universal hash function (in our case, SHA1 (sha1)), with the property that the probability of a hash collision is $1/N$ where $N$ is the range of possible hash values. Then, we can hash a vector $\bar{x}$ of $r$ integers as follows (carter1977universal):
@@
This approach enables us to map each band of MinHash signatures to a single integer value with known collision probability ($1/N$). To insert a document into the index, we simply insert each of the $b$ final integer values into each of the $b$ Bloom filters corresponding to the $b$ bands in the signature matrix.
4.3. Querying
Querying a document against the index is a similar process to insertion. We compute the document’s MinHash signature, group the rows into bands, and then hash each band as described in subsection 4.2. If there is a collision for any band of signatures, we mark that document as a duplicate. This approach mimics the original MinHashLSH procedure and gives the same probabilistic guarantees with respect to the Jaccard similarity of candidate pairs, with a slight increase in false positive probability due to our use of Bloom filters. However, the false positive probability is configurable and can be set to be arbitrarily low.
4.4. Error Rate
@@
Given a Minhash signature matrix with $b$ bands with $r$ rows each corresponding to $b$ Bloom filters with their individual false positive rates set to $p$, the net false positive overhead of using Bloom filters will be $p\textsubscript{effective}=1-(1-p)^{b}$. This is because a single collision in any band constitutes a duplicate, and false positive rates are independent for each Bloom filter. Therefore, the probability that no Bloom filter reports a false positive is $(1-p)^{b}$, and the overall false positive rate overhead introduced by using Bloom filters is the complement of this value.
@@
The relationship between the false positive rate of the Bloom filter and the overall error rate of LSHBloom can be understood as follows. leskovec2014finding determine that the false positive and false negative rate of MinHashLSH can be given by the following integrals, where $T$ is the Jaccard similarity threshold, $b$ is the number of bands in the MinHash signature matrix, and $r$ is the number of rows in each band:
(1) $FP_{lsh}=\int_{0}^{T}1-(1-t^{r})^{b}dt$
(2) $FN_{lsh}=\int_{T}^{1}1-(1-(1-t^{r})^{b})dt$
@@
zhu_lshensemble demonstrate how to tune the values of $b$ and $r$ to minimize these errors in MinhashLSH. We follow this approach in our work.
@@
A Bloom filter will not yield false negatives, but may yield false positives. If MinHashLSH yields a false positive (by “mistakenly” yielding two identical bands), we get a false positive result. If MinHashLSH does not yield a false positive, the bloom filter index may still yield a false positive result with probability $p_{effective}$. Additionally, there is a small probability of hash collision ($1/N$) when reducing each band to a single integer (as described in 4.2) for insertion into the Bloom filter. Therefore, the overall false positive rate for LSHBloom is given by:
(3) $FP_{bloom}=FP_{lsh}+(1-FP_{lsh})*(p_{effective}+b/N)$
Note that $p_{effective}$ can be set arbitrarily small in practice so that this overhead is negligible. For 32-bit hashvalues (as is the default in packages such as datasketch) $N=2^{32}$, so the probability of collision for bands prior to insertion into the Bloom filter is also quite small. Additionally, because the size of the Bloom filter index scales with the number of items inserted and is not influenced by the size of the integers inserted into it, one can increase the size of the hashvalues (to say, a 64 or 128-bit integer) without increasing the size of the Bloom filter index. This is not true of the traditional MinHashLSH index which must store these hashvalues for comparison.
Meanwhile, for each false negative that MinHashLSH produces, $p_{effective}$ percent of those false negatives will be marked as positives due to false positive hits in the bloom filter index and $b/N$ percent of them will be marked positive due to collisions when reducing the bands of the MinHash signature matrix for insertion. That is,
(4) $FN_{bloom}=(1-(p_{effective}+b/N))*FN_{lsh}$
4.5. Advantages
@@
We seek to measure duplication based on the Jaccard similarity between set-theoretic representations of documents as bags of n-grams. Since many practical scenarios may admit an extremely large universe of such n-grams (e.g., if we allow our set of documents to include a variety of encodings, languages, latex, tablature, figures, etc.), practitioners may want to reflect this additional complexity by increasing the size of MinHash hashvalues (to 64 or 128-bits or more) or increasing the accuracy of MinHash estimates of the Jaccard similarity by increasing the number of MinHash permutations. A scalable deduplication algorithm should be amenable to such considerations.
LSHBloom allows for the space-efficient representation of extremely large deduplication indices. This is due to the Bloom filter’s ability to determine the set membership properties of objects of arbitrary size with a fixed number of bits. This means that increasing the size of hashvalues does not change the index size. Furthermore, using more MinHash permutations only increases the index size linearly with the number of bands that are required for the MinHash signature matrix–which is far fewer than the number of permutations. If we expect $n$ documents and desire a false positive probability of $p$, then each Bloom filter requires exactly $m=-n*log(p)/(log(2))^{2}$ bits (bender2022optimal).
@@
As an example, say we use a Jaccard similarity threshold $T$ of 0.8, and 128 random permutations for MinHashing. For these parameter settings, MinHashLSH creates nine bands (using the method described in (zhu_lshensemble) to tune $b$ and $r$) for the MinHash signature matrix, meaning that we require nine Bloom filters for our index. For these settings, setting a conservative ppeffective = 1e-10 for a dataset of 10B documents would only require 590 GB of space. Meanwhile, traditional MinHashLSH would require 46 TB (nearly $80\times$ more space) to store the equivalent LSHIndex—exemplifying the scalability of our technique.
@@
5. Evaluation
Broadly our evaluation considers two concerns important for large-scale deduplication: correctness and performance. As expressed above, both the correctness and performance of these methods are critical for their application to extremely large datasets. To assess correctness, we tune LSHBloom and other state-of-the-art deduplication methods on a balanced dataset of 25K PDFs containing labeled duplicates. We then use the best parameter settings for each algorithm and compare them on a larger labelled dataset of 50K PDFs. Without sufficient correctness, the system risks either under-identifying duplicates, leading to decreased downstream model quality and unnecessary training overhead; or over-identifying duplicates, which unnecessarily reduces the training dataset volume. For performance, we consider the storage utilization and runtime of each method on a larger collection of 39M PDFs, peS2o (peS2o). Without sufficient performance, it may not be possible to process large datasets on available hardware resources.
@@ 5.1.2. Baselines
We compare LSHBloom with state-of-the-art deduplication methods widely used in large-scale LLM training: MinHashLSH, Dolma, Dolma-ngram, CCNet, and DataComp-LM. We normalized all implementations to Python-only implementations to avoid performance differences due to implementation language. This also enabled us to normalize the Bloom filter implementation for those techniques which use them. Below, we describe any changes or configurations to the methods.
@@
MinhashLSH (broder1997resemblance) (lsh): We configure the traditional MinhashLSH algorithm as described in subsection 2.3. DOLMA: We apply both standard dolma and dolma-ngram methods to document-level deduplication. Specifically, for dolma we determine whether a document is duplicated by identifying duplicated paragraphs and comparing the percentage of document text duplicated to the overlap threshold, $T$. For dolma-ngram, we split each document into a set of n-grams with a given size, determine how many n-grams are duplicated according to the Bloom filter, and compare the percentage of n-grams that are duplicated within that document to the overlap threshold, $T$. Our implementation aims to faithfully extend Dolma’s two primary deduplication methods to document-level deduplication. CCNet originally performs deduplication at the paragraph level; however, here, we apply it at the document level by measuring the proportion of duplicated paragraphs in any given document against some tolerance threshold, $T\in$ [0, 1]. For a fair comparison, we compare documents to all other documents—just as MinhashLSH does. When evaluating memory usage, we apply all CCNet optimizations, such as the use of memory-efficient hash sets. Additionally, we augment CCNet with multiprocessing where appropriate (e.g., when computing hashes over each document) for a more appropriate comparison during our scaling tests. DataComp-LM: For our evaluation, we compare only with their document-level deduplication procedure.
@@
When considering the correctness of techniques that operate on n-grams and use Bloom filters, we must give an expected n-gram count in the corpus; choosing this count accurately is crucial to make a fair comparison. Because counting the number of n-grams of a given size in a corpus is far more computationally expensive than counting the number of documents (as is required for our technique), we compute an estimate of the n-gram count in a corpus. To do so we randomly sample $N$ = 1000 documents from the corpus, compute the mean n-gram count from our sample, and multiply the mean n-gram count by the number of total documents in that corpus to obtain an estimate of the number of n-grams in the entire corpus. We parameterize Bloom filters for n-gram techniques using this estimate. We apply an analogous procedure to obtain estimated paragraph counts for Dolma and CCNet.
5.1.3. Metrics
@@
Correctness: We require methods that can distinguish meaningfully between duplicate and non-duplicate documents. Here, we consider a ”positive” label to mean that a document is a duplicate of another document in the corpus. We want these methods to yield a few false positives (to retrain the maximal number of training examples), few false negatives to reduce the prevalence of duplication in our dataset, and to provide a reasonable combination of accuracy and performance. Particularly for large-scale deep learning, retaining the maximal number of training examples is highly desirable—therefore, we want few false positives. Similarly, allowing a large number of duplicates into the training corpus is unacceptable as this affects the model’s training efficiency and may increase the likelihood that our model develops degenerate behaviors such as memorization. To account for these considerations, we evaluate each deduplication technique based on its ability to mitigate both false positives and false negatives using its precision (the proportion of duplicate labels that are true duplicates), its recall (the proportion of duplicates in the dataset that we correctly identify), and its F1 score (the harmonic mean of precision and recall):
@@
5.1.4. Constructing the correctness evaluation dataset
@@
To permit the rigorous evaluation of de-duplication methods, we created a pipeline for generating labeled datasets of duplicate documents. We acquire scientific articles (scraped from HTML and parsed from PDF) from a variety of sources, spanning a diverse set of disciplines (mathematics, physics, chemistry, biology, economics, engineering, and medicine) as identified through keywords or classification algorithms based on metadata. Text content from HTML is directly extracted. We translated each PDF into a sequence of tokens by applying a text extraction or OCR tool—one of PyMupDF (PyMuPDF), Nougat (blecher2023nougat), or Tesseract (smith2007overview), each used with roughly the same frequency. We then can construct datasets where documents are duplicated, either by including multiple versions of the same document which were parsed using different text extraction tools and/or by including randomly truncated versions of that document. This reflects common forms of duplication that occur in real-world web scraping pipelines where the same document can be ingested via different text extraction tooling either due to differences in the origin of the document instance, the fact that the same document is often found in multiple formats (HTML, XML, PDF, etc.), and due to parsing errors which can abruptly skip or truncate portions of the document’s text.
We build a balanced dataset of 24,000 documents for use in tuning hyperparameters for each deduplication algorithm. We further evaluate each technique using its best hyperparameter settings on datasets of 50,000 documents with increasing levels of duplication, from 10% to 90% duplicate instances.
5.1.5. Choice of values for Tuning Parameters
We evaluate each technique over a range of evenly spaced parameter settings using a grid search. Specifically, for any threshold parameters such as the Jaccard similarity threshold for MinHashLSH and the text overlap threshold for DataComp-LM, we evaluate values between 0.2 and 1.0 in steps of 0.2. We vary the number of permutations for MinHashLSH from 32 to 256 by powers of two. For n-gram techniques, we explore n-gram sizes 1, 2, 5, 7, 13, and 26. For MinHashLSH and LSHBloom, we additionally gathered results for $T=0.5$ and 48 random permutations as the surrounding parameter values demonstrated sharp jumps in the F1 score, and we wanted to explore this change in the score at a finer granularity. All Bloom filters (for DOLMA/DOLMA-ngram, DataComp-LM, LSHBloom) are assigned a false positive rate $p$ of 1e-5. For LSHBloom, this equates to $p\textsubscript{effective}$ between 1e-5 and 8e-4, depending on the value of $T$ and the number of MinHash permutations. The value of $p_{effective}$ at various parameter settings for this benchmark is illustrated in Figure 2.